### v1

@@ 3.  Method
In this section, we provide a comprehensive overview of our method. Initially, we observe a pronounced transferability of gradient information around the clean images. Stemming from this observation, we introduce a novel concept termed Neighbourhood Gradient Information. To harness the full potential of this gradient information, we propose the NGI-Attack, incorporating two key strategies: Example Backtracking (EB) and Multiplex Mask (MM). In particular, EB allows us to accumulate such gradients via a momentum term. However, given the finite number of iterations, accumulating Neighborhood Gradient Information consumes valuable iteration time. Consequently, increasing the iteration step size is crucial to reaching maximum perturbation magnitude within fewer iterations. Further, to induce significant perturbation without overfitting to the attack, we apply the MM strategy to form a multi-way attack. A more detailed description of our methods follows.
@@
Figure 1: Two toy experiments using Inception-v3 (Szegedy et al., 2016) as the surrogate model: (a) examines the impact of accumulating Neighborhood Gradient Information, and (b) validates the effects of direct amplification steps on attack performance.
@@ 3.1. Preliminaries
Denote $f$: $x$rightarrow $y$ as the classification process of a surrogate model, where $x$ is the clean image and $y$ is the corresponding ground-truth label. Let $\mathcal{J}(x,y)$ represent the loss function (e.g. the cross-entropy loss). To obtain adversarial examples $x^{adv}$, we need to add a perturbation to $x$ to maximize the loss $\mathcal{J}(x^{adv},y)$ such that the target model $f$ makes an incorrect classification prediction, i.e. $f(x^{adv})\neq y$ (non-targeted attack). Following  (Xie et al., 2019; Dong et al., 2019; Lin et al., 2019; Dong et al., 2017; Wang & He, 2021), we adopt the $L_{\infty}$ norm, whereby $\|x-x^{adv}\|_{\infty}\leq\epsilon$, with $\epsilon$ serving as the upper bound on the constrained perturbation to ensure attack imperceptibility. In the black-box transfer attack, where direct access to the target model $f$ is not available, a surrogate model must be identified as a target for the attack. So the optimization objective can be summarised as:
$\mathop{\mathrm{argmax}}\limits_{x_{adv}}(\mathcal{J}(x^{adv},y)),\quad s.t.~{}||x-x^{adv}||_{\infty}\leq\epsilon.$
Goodfellow et al. (2015) first propose the Fast Gradient Sign Method (FGSM), which generates adversarial examples through a single-step attack. Subsequently Kurakin et al. (2018) propose the Iterative Fast Gradient Sign Method (I-FGSM), and most gradient optimization-based transfer attack methods are based on I-FGSM. With a total of $T$ iterations, constrain each attack step to be $\alpha=\epsilon/T$. The attack process can be described as follows:
$x^{adv}_{t+1}=Clip_{\epsilon}(x^{adv}_{t}+\alpha\cdot{\rm sign}(\nabla_{x}\mathcal{J}(x^{adv}_{t},y))),$
@@
Figure 2: Visualization of the attention of adversarial examples under different strategies with Grad-cam. The top row, the middle row, and the bottom row represent the attention distribution of the original image, the generated adversarial examples under the scale-up step strategy, and the generated adversarial examples under our strategy, respectively, under different defense models.
@@ 3.2. Motivation
To further verify whether the Neighbourhood Gradient Information enhances the attack transferability, we conduct an experiment. Specifically, we introduce additional rounds of attacks near the clean images before executing attack, and we use momentum to accumulate the Neighbourhood Gradient Information. As shown in Figure 1a, collecting the Neighbourhood Gradient Information for just one round in advance significantly improves the attack success rate and reaches a high attack success rate at 5 rounds of collection. However, as the number of iterations increases, the attack effectiveness tends to stabilize. This indicates that collecting gradient information from later iterations, when the adversarial examples are away from the clean images, has a diminishing impact on improving adversarial transferability. In conclusion, the perturbations generated around the clean images are indeed more effective for the attack.
@@
To maximize the utilization of the Neighbourhood Gradient Information without additional time costs, we attempt to increase the step size in each iteration, thereby enhancing the perturbation magnitude. However, we observe that directly increasing the step size may weaken the transferability of adversarial examples. As illustrated in Figure 1b, when the step size is doubled, i.e., step-scaling, there is a slight increase in the attack success rate on normally trained models, highlighting the effectiveness of Neighbourhood Gradient Information. However, against adversarially trained models, there is a notable decrease, ranging from 5.2% to 9.3%. We infer that excessively large attack steps may lead to overfitting of the attack.To further elucidate the underlying factors contributing to this phenomenon, we employ Grad-CAM (Selvaraju et al., 2020) to visually analyze the behavior of these defense models. Based on the observation depicted in Figure 2, where Grad-CAM refocuses the region of interest associated with the original network classification when the step size is directly scaled up, we infer that excessively large attack steps may lead to overfitting of the attack. Therefore, the question arises: Can we find a more straightforward and effective method that improves the attack success rate without incurring the additional time cost associated with accumulating the Neighbourhood Gradient Information?
@@ 3.3. NGI-Attack
To address the aforementioned issue and effectively leverage the Neighbourhood Gradient Information, we propose the NGI-Attack, incorporating two key strategies: Example Backtracking (EB) and Multiplex Mask (MM).
@@
Example Backtracking. EB involves a direct accumulation of Neighbourhood Gradient Information, which is subsequently utilized to generate the adversarial examples. Moreover, to enhance the collection and utilization of gradient information near the clean images, additional processing of the obtained gradient information is necessary when employing I-FGSM. As a result, we directly utilize MI-FGSM as our baseline, which incorporates momentum to collect the gradient information effectively. To mitigate the introduction of additional time costs, we adopt a two-stage approach in the overall attack process. Specifically, during the initial $K$ rounds, we focus on adjacent gradient collection, while in the remaining $(T-K)$ rounds, we conduct substantial attacks to generate adversarial examples. Thus, the gradient collection process during the attack is as follows:
$\begin{split}g_{t+1}&=\mu\cdot g_{t}+\frac{\nabla_{x}\mathcal{J}(x^{adv}_{t},y)}{||\nabla_{x}\mathcal{J}(x^{adv}_{t},y)||_{1}},\\ x^{adv}_{t+1}&=\text{Clip}_{\epsilon}\left(x^{adv}_{t}+\alpha\cdot\text{sign}(g_{t+1})\right),\end{split}$ (1)
@@
where $g_{t}$ gathers gradients of previous $t$ iterations with an initial value of $g_{0}$ = 0 and $\mu$ is the decay factor of momentum term. The Example Backtracking is also easy to perform, simply resetting the example back to the clean image when the number of iteration rounds reaches $K$. That is, when the number of iteration rounds $t$ in the iterative attack is equal to $K$, reset $x^{adv}_{K}$ = $x^{adv}_{0}$. But direct use of Example Backtracking inevitably poses a problem: to ensure that the attack reaches the maximum perturbation limit without extra time costs, it is necessary to employ a larger step size in the latter $(T-K)$ rounds. However, as indicated by the previous analysis, directly increasing the step size may lead to overfitting of the attack. Therefore, obtaining more effective information about the attack gradient within the same iteration rounds becomes a crucial factor in improving the attack success rate.
@@
Figure 3: Visualization of the attention of MaskProcess images with Grad-cam towards Inception-v3 (Szegedy et al., 2016). Each set comprises two rows; the first represents the original image and its attention visualization, while the second represents the MaskProcess image with $P$ of 0.9 and its attention visualization.
@@
Multiplex Mask. Direct use of increasing step size leads to overfitting of the attack process. We attribute this phenomenon to the repetitive use of the same gradient information. A more sophisticated approach is to gather diverse gradient information. This allows us to effectively split the larger step size, rather than linearly scaling it, which in turn suppresses the aforementioned overfitting issue. The defense model turns out to have a more expansive attention region for the target compared to the normally trained model (Dong et al., 2019), and according to Figure 1b, when images are subjected to attacks with excessively large step sizes, they may weaken transferability towards adversarially trained models. To counter this, we propose to augment the input with a mask. This makes the model focus on more non-discriminative regions, effectively expanding the attention region. Subsequently, it enables a more comprehensive gradient information extraction and a more thorough feature destruction.
@@
To visually underscore this phenomenon, we present the MaskProcess in Figure 3. From the figure, it’s evident how the network’s attention shifts towards the non-discriminative regions after applying the masking strategy. By using the masked-enhanced image, we can obtain new gradient information and capture the gradients of the attack on different feature regions:
$\begin{split}x_{t}^{mask}&=\operatorname{MaskProcess}(x_{t}^{adv},P),\\ g_{t+1}^{mask}&=\mu\cdot g_{t}^{mask}+\frac{\nabla_{x}\mathcal{J}(x^{mask}_{t},y)}{||\nabla_{x}\mathcal{J}(x^{mask}_{t},y)||_{1}},\end{split}$ (2)
@@
Time Complexity Analysis. Our NGI-Attack method enhances adversarial transferability without introducing additional time costs. During the first $K$ iterations, we apply the EB strategy to gather Neighbourhood Gradient Information. For the subsequent $(T-K)$ iterations, we employ the MM strategy, a multi-way attack approach where it processes two inputs simultaneously. These operations are completed within an $O(1)$ time complexity and do not incur additional computational overhead. Therefore, despite involving more complex operations, the overall time cost of our method does not increase, maintaining the total number of iterations at $T$. This aligns with the time complexity of previous iterative attack methods  (Dong et al., 2018; Xie et al., 2019; Dong et al., 2019).
@@ 4.  Experiments
To validate the effectiveness of our proposed NGI-Attack, we conduct comprehensive experiments on the standard ImageNet dataset (Deng et al., 2009). This section initiates with a thorough outline of our experimental setup. Subsequently, we compare our approach with competitive baselines across varied experimental conditions and provide a quantitative analysis of attack performance on nine advanced defense models. Results consistently underscore that our method significantly enhances transferability across diverse scenarios. For further investigation, we also examine the effects of different proposed modules and the influence of hyperparameters, namely the Example Backtracking step $K$ and the Multiplex Mask probability $P$.
@@ 4.1. Experiment Setup
Datasets. Following (Kurakin et al., 2017; Wang & He, 2021; Long et al., 2022; Dong et al., 2017; Lin et al., 2019), we utilize the dataset employed in the NIPS 2017 adversarial competition [^1] https://github.com/cleverhans-lab/cleverhans/tree/master/cleverhans˙v3.1.0/examples/nips17˙adversarial˙competition/dataset, consisting of 1000 images randomly picked from the ILSVRC 2012 validation set. These images are ensured to remain within the 1000 categories and to be accurately classified by the test models.
Models. In our study, we undertake a comprehensive evaluation of diverse models. We consider four normally trained models, comprising Inception-v3 (Inc-v3) (Szegedy et al., 2016), Inception-v4 (Inc-v4) (Szegedy et al., 2017), Inception-Resnet-v2 (IncRes-v2), and Resnet-v2-101 (Res-101) (He et al., 2016). Additionally, we evaluate two recent transformer-based models, namely Vision Transformer (ViT) (Dosovitskiy et al., 2021) and Swin Transformer (Swin-b) (Liu et al., 2021). Furthermore, we assess three adversarially trained models, which are ens3-adv-Inception-v3 (Inc-v3ens3), ens4-adv-Inception-v3 (Inc-v3ens4) and ens-adv-Inception-ResNet-v2 (IncRes-v2ens(Tramèr et al., 2018). Beyond the models, our study also delves into nine advanced defense methods renowned for their robustness against black-box attacks on the ImageNet dataset, specifically HGD (Liao et al., 2018), R&P (Xie et al., 2017), NIPS-r3222https://github.com/anlthms/nips-2017/tree/master/mmd, Bit-Red (Xu et al., 2018), JPEG (Guo et al., 2017), FD (Liu et al., 2019), ComDefend (Jia et al., 2019), NRP (Naseer et al., 2020), and RS (Cohen et al., 2019). We apply the official models for the first three, while for the remaining models, we test their effectiveness using Inc-v3ens3.
Baselines. We select two typical optimization-based attacks, specifically MI-FGSM (Dong et al., 2018) and NI-FGSM (Lin et al., 2020), as our baseline comparisons. For input transformation-based attacks, we opt for DI-FGSM (Xie et al., 2019), TI-FGSM (Dong et al., 2019), and SI-FGSM  (Lin et al., 2019) as our baselines. Moreover, to validate the efficacy of our approach, we contrast it with various recent state-of-the-art attack methods, including SI-NI-FGSM, VT-FGSM (Wang & He, 2021), Admix (Wang et al., 2021), SSA (Long et al., 2022), RAP (Qin et al., 2022), SIA (Wang et al., 2023), BSR (Wang et al., 2024), and combined variants of these methods, such as VT-TI-DIM (combined version VT-FGSM, DI-FGSM, and TI-FGSM).
@@
Table 1: Attack success rates (%) of optimization-based attacks on nine normally and adversarially trained models in the single-model setting. The adversarial examples are crafted on Inc-v3, Inc-v4, IncRes-v2, and Res101 respectively. * indicates the white-box model.

### v2

@@ 3. Method
In this section, we provide a comprehensive overview of our method. Initially, we observe a pronounced transferability of gradient information around the clean images. Stemming from this observation, we introduce a novel concept termed Neighbourhood Gradient Information. To harness the full potential of this gradient information, we propose the NGI-Attack, incorporating two key strategies: Example Backtracking (EB) and Multiplex Mask (MM). In particular, EB allows us to accumulate such gradients via a momentum term. However, given the finite number of iterations, accumulating Neighborhood Gradient Information consumes valuable iteration time. Consequently, increasing the iteration step size is crucial to reaching maximum perturbation magnitude within fewer iterations. Further, to induce significant perturbation without overfitting to the attack, we apply the MM strategy to form a multi-way attack.
@@
Figure 1: Two toy experiments using Inception-v3 (inceptionv2) as the surrogate model: (a) examines the impact of accumulating Neighborhood Gradient Information, and (b) validates the effects of direct amplification steps on attack performance.
@@ 3.1. Preliminaries
Denote $f$: $x$ $\rightarrow$ $y$ as the classification process of a surrogate model, where $x$ is the clean image and $y$ is the corresponding ground-truth label. Let $\mathcal{J}(x,y)$ represent the loss function (e.g. the cross-entropy loss). To obtain adversarial examples $x^{adv}$, we need to add a perturbation to $x$ to maximize the loss $\mathcal{J}(x^{adv},y)$ such that the target model $f$ makes an incorrect classification prediction, i.e. $f(x^{adv})\neq y$ (non-targeted attack). Following  (xie2019improving; dong2019evading; DBLP:journals/corr/abs-1908-06281; dong2017discovering; DBLP:conf/cvpr/Wang021), we adopt the $L_{\infty}$ norm, whereby $\|x-x^{adv}\|_{\infty}\leq\epsilon$, with $\epsilon$ serving as the upper bound on the constrained perturbation to ensure attack imperceptibility. In the black-box transfer attack, where direct access to the target model $f$ is not available, a surrogate model must be identified as a target for the attack. So the optimization objective can be summarised as:
[…]
fgsm first propose the Fast Gradient Sign Method (FGSM), which generates adversarial examples through a single-step attack. Subsequently ifgsm propose the Iterative Fast Gradient Sign Method (I-FGSM), and most gradient optimization-based transfer attack methods are based on I-FGSM. With a total of $T$ iterations, constrain each attack step to be $\alpha=\epsilon/T$. The attack process can be described as follows:
$x^{adv}_{t+1}=Clip_{\epsilon}(x^{adv}_{t}+\alpha\cdot{\rm sign}(\nabla_{x}\mathcal{J}(x^{adv}_{t},y))),$ (10)
@@
Figure 2: Visualization of the attention of adversarial examples under different strategies with Grad-Cam. The top row, the middle row, and the bottom row represent the attention distribution of the clean image, the generated adversarial examples under the scale-up step strategy, and the generated adversarial examples under our strategy, respectively, under different defense models. The results demonstrate that our strategy successfully achieves adversarial attacks by effectively disrupting the network’s attention, while the scale-up strategy tends to maintain regions of interest similar to those in the clean images across various defense models.
@@ 3.2. Motivation
To further verify whether the Neighbourhood Gradient Information enhances the attack transferability, we conduct an experiment. Specifically, we introduce additional rounds of attacks near the clean images before executing attack, and we use momentum to accumulate the Neighbourhood Gradient Information. As shown in Figure 1a, collecting the Neighbourhood Gradient Information for just one round in advance significantly improves the attack success rate and reaches a high attack success rate at 5 rounds of collection. However, as the number of iterations increases, the attack effectiveness tends to stabilize. This indicates that collecting gradient information from later iterations, when the adversarial examples are away from the clean images, has a diminishing impact on improving adversarial transferability. In conclusion, the perturbations generated around the clean images are indeed more effective for the attack. This empirical finding is consistent with recent theoretical analyses showing that early-stage gradients reside in low-frequency, flat regions of the loss landscape—features widely shared across models and thus highly transferable (pmlr-v97-rahaman19a; sagun2017eigenvalues; cohen2021gradient; foret2021sharpnessaware; Yinrobustness; Lidefense; hu2020adversarial).
@@
To maximize the utilization of the Neighbourhood Gradient Information without additional time costs, we attempt to increase the step size in each iteration, thereby enhancing the perturbation magnitude. However, we observe that directly increasing the step size may weaken the transferability of adversarial examples. As illustrated in Figure 1b, when the step size is doubled, i.e., step-scaling, there is a slight increase in the attack success rate on normally trained models, highlighting the effectiveness of Neighbourhood Gradient Information. However, against adversarially trained models, there is a notable decrease, ranging from 5.2% to 9.3%. To further elucidate the factors underlying this phenomenon, we employ Grad-Cam (gradcam) to analyze the behavior of various defense models visually. Figure 2 presents the clean images, adversarial examples generated using the scale-up step strategy, and those produced by our approach, along with their corresponding Grad-Cam heatmaps across different defense models. The results indicate that successful adversarial attacks disrupt the network’s attention, as evidenced by significant divergence in Grad-Cam heatmaps from the clean images. Conversely, unsuccessful attacks tend to maintain regions of interest similar to those in the clean images. Our findings reveal that adversarial examples generated by the scale-up strategy exhibit consistent attention patterns across defense models, indicating limited transferability, likely due to the overfitting associated with excessively large attack steps. In contrast, our method effectively disrupts attention across various defense models, thereby demonstrating enhanced transferability. Therefore, the question arises: Can we find a more straightforward and effective method that improves the attack success rate without incurring the additional time cost associated with accumulating the Neighbourhood Gradient Information?
Figure 3: The overall NGI-Attack pipeline: The process begins using the Example Backtracking strategy to collect Neighbourhood Gradient Information. This information is then utilized in the Multiplex Mask strategy, forming a multi-way attack. One pathway applies the MaskProcess operation, while the other directly utilizes the clean image. The gradients of these two pathways are collected independently, guided by the Neighbourhood Gradient Information, ultimately generating a highly transferable adversarial example.
@@ 3.3. NGI-Attack
To address the aforementioned issue and effectively leverage the Neighbourhood Gradient Information, we propose the NGI-Attack, incorporating two key strategies: Example Backtracking (EB) and Multiplex Mask (MM). The overall NGI-Attack pipeline is depicted in Figure 3.
@@
Example Backtracking. EB performs a Neighbourhood Gradient Collection, directly accumulating gradient information, which is subsequently utilized to generate adversarial examples. Moreover, to enhance the collection and utilization of Neighbourhood Gradient Information, additional processing of the obtained gradient information is necessary when employing I-FGSM. As a result, we directly utilize MI-FGSM as our baseline, which incorporates momentum to collect the gradient information effectively. To mitigate the introduction of additional time costs, we adopt a two-stage approach in the overall attack process. Specifically, we dedicate the initial $K$ rounds to Neighbourhood Gradient Collection, followed by the remaining (TK)(T-K) rounds, where we conduct substantial attacks to generate adversarial examples. Thus, the Neighbourhood Gradient Collection process during the attack is as follows:
[…]
where $g_{t}$ gathers gradients of previous $t$ iterations with an initial value of $g_{0}$ = 0 and $\mu$ is the decay factor of momentum term. Subsequently, Example Backtracking performs a simple reset of the example back to the clean image after finishing $K$ rounds of Neighbourhood Gradient Collection. That is, when the number of iteration rounds $t$ in the iterative attack is equal to $K$, reset $x^{adv}_{K}$ = $x^{adv}_{0}$. To formalize the gradient accumulation during this phase, we define the Neighbourhood Gradient Information (NGI) as:
$g_{\text{NGI}}=g_{K}=\sum_{t=1}^{K}\mu^{K-t+1}\cdot\frac{\nabla_{x}\mathcal{J}(x_{t-1}^{adv},y)}{\left\|\nabla_{x}\mathcal{J}(x_{t-1}^{adv},y)\right\|_{1}}.$ (12)
@@
This $g_{\text{NGI}}$ then serves as the initialized momentum term for the subsequent (TK)(T-K) attack rounds. But direct use of Example Backtracking inevitably poses a problem: to ensure that the attack reaches the maximum perturbation limit without extra time costs, it is necessary to employ a larger step size in the latter (TK)(T-K) rounds. However, as indicated by the previous analysis, directly increasing the step size may lead to overfitting of the attack. Therefore, obtaining more effective information about the attack gradient within the same iteration rounds becomes a crucial factor in improving the attack success rate.
@@
Figure 4: Visualization of the attention of MaskProcess images with towards Inception-v3 (inceptionv2). Each set comprises two rows; the first represents the clean image and its attention visualization, while the second represents the MaskProcess image with $P$ of 0.9 and its attention visualization. The results show that the MaskProcess image can successfully shift the network’s attention towards non-discriminative regions.
@@
Multiplex Mask. Direct use of increasing step size leads to overfitting of the attack process. We attribute this phenomenon to the repetitive use of the same gradient information. A more sophisticated approach is to gather diverse gradient information. This allows us to effectively split the larger step size, rather than linearly scaling it, which in turn suppresses the aforementioned overfitting issue. The defense model turns out to have a more expansive attention region for the target compared to the normally trained model (dong2019evading), and according to Figure 1b, when images are subjected to attacks with excessively large step sizes, they may weaken transferability towards adversarially trained models. To counter this, we propose to augment the input with a mask. This makes the model focus on more non-discriminative regions, effectively expanding the attention region. Subsequently, it enables a more comprehensive gradient information extraction and a more thorough feature destruction.
@@
To visually underscore this phenomenon, we present the MaskProcess in Figure 4. The figure shows that the network’s attention shifts towards the non-discriminative regions after applying the masking strategy. By using the mask-enhanced image, we capture new gradient information across different feature regions, thereby enhancing the adversarial transferability:
[…]
@@
Time Complexity Analysis. Our NGI-Attack method enhances adversarial transferability without introducing additional time costs. During the first $K$ iterations, we apply the EB strategy to gather Neighbourhood Gradient Information. For the subsequent (TK)(T-K) iterations, we employ the MM strategy, a multi-way attack approach that processes two inputs simultaneously. These operations are completed within an $O(1)$ time complexity and do not incur additional computational overhead. Therefore, despite involving more complex operations, the overall time cost of our method does not increase, maintaining the total number of iterations at $T$. This aligns with the time complexity of previous iterative attack methods  (mi; xie2019improving; dong2019evading).
To further validate the theoretical analysis, we conduct an empirical runtime evaluation using Inception-v3 as the surrogate model. All methods are implemented under the same settings (batch size = 10) on an NVIDIA Tesla V100 (32GB). The results show that NGI-Attack only incurs a moderate increase in runtime compared to MI-FGSM and PGD (1.38s vs. 0.87s/0.88s per batch), while being significantly faster than VMI-FGSM (11.63s per batch). These results confirm that our method improves adversarial transferability without introducing noticeable runtime overhead.
@@ 4. Experiments
To validate the effectiveness of our proposed NGI-Attack, we conduct comprehensive experiments on the standard ImageNet dataset (imagenet). This section initiates with a thorough outline of our experimental setup. Subsequently, we compare our approach with competitive baselines across varied experimental conditions and provide a quantitative analysis of attack performance on nine advanced defense models. Results consistently underscore that our method significantly enhances transferability across diverse scenarios. For further investigation, we also examine the effects of different proposed modules and the influence of hyperparameters, namely the Example Backtracking step $K$ and the Multiplex Mask probability $P$.
@@ 4.1. Experiment Setup
Datasets. Following (DBLP:conf/iclr/KurakinGB17; DBLP:conf/cvpr/Wang021; ssa; dong2017discovering; DBLP:journals/corr/abs-1908-06281), we utilize the dataset employed in the NIPS 2017 adversarial competition [^1] https://github.com/cleverhans-lab/cleverhans/tree/master/cleverhans˙v3.1.0/examples/nips17˙adversarial˙competition/dataset, consisting of 1000 images randomly picked from the ILSVRC 2012 validation set. These images are ensured to remain within the 1000 categories and to be accurately classified by the test models.
Models. In our study, we undertake a comprehensive evaluation of diverse models. We consider four normally trained models, comprising Inception-v3 (Inc-v3) (inceptionv2), Inception-v4 (Inc-v4) (szegedy2017inception), Inception-Resnet-v2 (IncRes-v2), and Resnet-v2-101 (Res-101) (DBLP:conf/eccv/HeZRS16). Additionally, we evaluate two recent transformer-based models, namely Vision Transformer (ViT) (vit) and Swin Transformer (Swin-b) (swint). Furthermore, we assess three adversarially trained models, which are ens3-adv-Inception-v3 (Inc-v3ens3), ens4-adv-Inception-v3 (Inc-v3ens4) and ens-adv-Inception-ResNet-v2 (IncRes-v2ens(Florian2018Ensemble). Beyond the models, our study also delves into nine advanced defense methods renowned for their robustness against black-box attacks on the ImageNet dataset, specifically HGD (DBLP:conf/cvpr/LiaoLDPH018), R&P (xie2017mitigating), NIPS-r3222https://github.com/anlthms/nips-2017/tree/master/mmd, Bit-Red (DBLP:conf/ndss/Xu0Q18), JPEG (guo2017countering), FD (liu2019feature), ComDefend (DBLP:conf/cvpr/JiaWCF19), NRP (DBLP:conf/cvpr/NaseerKHKP20), and RS (DBLP:conf/icml/CohenRK19). We apply the official models for the first three, while for the remaining models, we test their effectiveness using Inc-v3ens3.
Baselines. We select three typical optimization-based attacks, specifically MI-FGSM (mi), NI-FGSM (sini), and VMI-FGSM (DBLP:conf/cvpr/Wang021), as our baseline comparisons. For input transformation-based attacks, we opt for DI-FGSM (xie2019improving), TI-FGSM (dong2019evading), and SI-FGSM  (DBLP:journals/corr/abs-1908-06281) as our baselines. We also include SGM (sgm) as a baseline. Moreover, to validate the efficacy of our approach, we contrast it with various recent state-of-the-art attack methods, including SI-NI-FGSM, VT-FGSM (DBLP:conf/cvpr/Wang021), Admix (wang2021admix), SSA (ssa), RAP (NEURIPS2022_c0f9419c), SIA (SIA), BSR (wang2024boosting), and combined variants of these methods, such as VT-TI-DIM (a combined version of VT-FGSM, DI-FGSM, and TI-FGSM).
@@
Table 1: Attack success rates (%) of optimization-based attacks on nine normally and adversarially trained models in the single-model setting. The adversarial examples are crafted on Inc-v3, Inc-v4, IncRes-v2, and Res101, respectively. * indicates the white-box model.